OpenStudy (anonymous):

Linda would like to hire a clown to perform at her son’s birthday party. Clown City charges $100 for incidentals plus $30 per hour for hiring a clown to perform. Party Performers charges $120 for incidentals plus $25 per hour. How many hours would Linda need to plan on so that Party Performers is less expensive than Clown City?

OpenStudy (anonymous):

We need to equations. For the first company I have y = 100 + 30x Can you do the same for the other company?

OpenStudy (anonymous):

like this y = 125 + 25x

OpenStudy (anonymous):

Yes, now we can set them equal and solve for x 100+30x = 125 + 25x What is x?

OpenStudy (anonymous):

I think I am suppose to use the > sign as well

OpenStudy (anonymous):

Opps, did you mean 120 not 125? We can change the equal sign 100+30x > 120 + 25x What is x?

OpenStudy (anonymous):

yes sorry lol

OpenStudy (anonymous):

thanks a bunch!

OpenStudy (anonymous):

x > 4

OpenStudy (anonymous):

Correct. So how many hours would Linda need to plan on so that Party Performers is less expensive than Clown City?

OpenStudy (anonymous):

she would need to plan the Party Performers for more than 4 hours